(a) (a) What energy (in J) is dissipated by a lightning bolt having a 17,000 A...

(a)

(a) What energy (in J) is dissipated by a lightning bolt having a 17,000 A current, a voltage of 1.45 ✕ 102 MV, and a length of 1.00 ms?

___________J

(b) What mass (in kg) of tree sap could be raised from 18.0°C to its boiling point and then evaporated by this energy, assuming sap has the same thermal characteristics as water?

___________kg
Answer #1

© Evit E = mit = 9oys X10 2 X106 x 17000 x 1410-3 E = 2.514X109 J E = mSAT + m2 2.514x109 m 14184) (100-18) + m (2256X10) Tm=